# Doubly Optimal Policy Evaluation for Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement learning (RL, Sutton and Barto, (2018)) has achieved remarkable success in various sequential decision-making problems. For example, RL algorithms have reduced energy consumption for Google data centers’ cooling by $40\%$ (Chervonyi et al.,, 2022), predicted protein structures with competitive accuracy (Jumper et al.,, 2021), and discovered faster matrix multiplication algorithms (Fawzi et al.,, 2022). When applying RL algorithms, policy evaluation plays a critical role to allow practitioners to estimate the performance of a policy before committing to its full deployment and test different algorithmic choices. A commonly used approach among RL practitioners for policy evaluation is the on-policy Monte Carlo method, where a policy (i.e., the target policy) is evaluated by directly executing itself. However, using the target policy itself as the behavior policy is not optimal (Hanna et al.,, 2017; Liu and Zhang,, 2024), leading to evaluation with high variance. This suboptimality of on-policy evaluation results in extensive needs for collecting online samples to achieve a desired level of accuracy.
@@ 2 Related Work
Reducing the variance for policy evaluation in reinforcement learning (RL) has been widely studied. One rising approach is variance reduction by designing a proper data-collecting policy, also known as the behavior policy. Noticing that the target policy itself is not the best behavior policy, Hanna et al., (2017) formulate the task of finding a variance-reduction behavior policy as an optimization problem. They use stochastic gradient descent to update a parameterized behavior policy. However, the stochastic method has been known to easily get stuck in highly suboptimal points in just moderately complex environments, where various local optimal points exist (Williams,, 1992). By contrast, our method directly learns the globally optimal behavior policy without doing a policy search. Moreover, their method requires highly sensitive hyperparameter tuning to learn the behavior policy effectively. Specifically, the learning rate can vary by up to $10^{5}$ times across different environments, as reported in the experiments of Hanna et al., (2017). This extreme sensitivity requires online tuning, consuming massive online data. By contrast, we propose an efficient algorithm to learn our optimal behavior policy with purely offline data. Furthermore, Hanna et al., (2017) constrain the online data to be complete trajectories. By contrast, our method copes well with incomplete offline tuples, which is widely applicable.
@@ 3 Background
In this paper, we study a finite horizon Markov Decision Process (MDP, Puterman, (2014)). This MDP consists of a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition probability function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1]$, and a constant horizon length $T$. For simplifying notations. we consider the undiscounted setting without loss of generality. Our method naturally applies to the discounted setting as long as the horizon is fixed and finite (Puterman,, 2014). We define a shorthand $[n]\doteq\quantity{0,1,\dots,n}$ for any integer $n$.
@@
The MDP process begins at time step 00, where an initial state $S_{0}$ is sampled from $p_{0}$. At each time step $t\in[T-1]$, an action $A_{t}$ is sampled based on $\pi_{t}(\cdot\mid S_{t})$. Here, $\pi_{t}:\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the policy at time step $t$. Then, a finite reward $R_{t+1}\doteq r(S_{t},A_{t})$ is emitted and a successor state $S_{t+1}$ is sampled from $p(\cdot\mid S_{t},A_{t})$. We also define abbreviations $\pi_{i:j}\doteq\quantity{\pi_{i},\pi_{i+1},\dots,\pi_{j}}$ and $\pi\doteq\pi_{0:T-1}$. At each time step $t$, the return is defined as $G_{t}\doteq\sum_{i={t+1}}^{T}R_{i},$ which allows us to define the state-value and action-value functions as $v_{\pi,t}(s)\doteq\mathbb{E}_{\pi}\left[G_{t}\mid S_{t}=s\right]$ and $q_{\pi,t}(s,a)\doteq\mathbb{E}_{\pi}\left[G_{t}\mid S_{t}=s,A_{t}=a\right].$ We use the total rewards performance metric (Puterman,, 2014) to measure the performance of the policy $\pi$, which is defined as $J(\pi)\doteq\sum_{s}p_{0}(s)v_{\pi,0}(s)$. In this paper, we use Monte Carlo methods, as introduced by Kakutani, (1945), for estimating the total rewards $J(\pi)$. The most straightforward Monte Carlo method is to draw samples of $J(\pi)$ through the online execution of the policy $\pi$. The empirical average of the sampled returns converges to $J(\pi)$ as the number of samples increases. Since this method estimates a policy $\pi$ by executing itself, it is called on-policy learning (Sutton, 1988).
@@
Moving forward, we focus on off-policy evaluation, where the goal is to estimate the total rewards $J(\pi)$ of an interested policy $\pi$, which is called the target policy. Data for off-policy evaluation are collected by executing a different policy $\mu$, called the behavior policy. In off-policy evaluation, we generate each trajectory $\quantity{S_{0},A_{0},R_{1},S_{1},A_{1},R_{2},\dots,S_{T-1},A_{T-1},R_{T}}$ by a behavior policy $\mu$ with $A_{t}\sim\mu_{t}(\cdot|S_{t}).$ We use a shorthand $\tau^{\mu_{t:T-1}}_{t:T-1}\doteq\quantity{S_{t},A_{t},R_{t+1},\dots,S_{T-1},A_{T-1},R_{T}}$ for a trajectory generated by the behavior policy $\mu$ from the time step $t$ to the time step $T-1$ inclusively. In off-policy learning, we use the importance sampling ratio to reweight rewards collected by the behavior policy $\mu$, in order to give an estimate of $J(\pi)$. We define the importance sampling ratio at time step $t$ as $\rho_{t}\doteq\frac{\pi_{t}(A_{t}\mid S_{t})}{\mu_{t}(A_{t}\mid S_{t})}.$ The product of importance sampling ratios from time $t$ to $t^{\prime}\geq t$ is then defined as $\rho_{t:t^{\prime}}\doteq\prod_{k=t}^{t^{\prime}}\frac{\pi_{k}(A_{k}|S_{k})}{\mu_{k}(A_{k}|S_{k})}.$ There are several ways to use the importance sampling ratios in off-policy learning (Geweke,, 1988; Hesterberg,, 1995; Koller and Friedman,, 2009; Thomas,, 2015). In this paper, we investigate the per-decision importance sampling estimator (PDIS, Precup et al., (2000)) and leave the investigation of others for future work. The PDIS Monte Carlo estimator is defined as $G^{\text{PDIS}}(\tau^{\mu_{t:T-1}}_{t:T-1})\doteq\sum_{k=t}^{T-1}\rho_{t:k}R_{k+1},$ which can also be expressed recursively as
@@
Under the classic policy coverage assumption (Precup et al.,, 2000; Maei,, 2011; Sutton et al.,, 2016; Zhang,, 2022) $\forall t,s,a,\mu_{t}(a|s)=0\implies\pi_{t}(a|s)=0$, this off-policy estimator $G^{\text{PDIS}}(\tau^{\mu_{0:T-1}}_{0:T-1})$ provides an unbiased estimation for $J(\pi)$, i.e., $\mathbb{E}\quantity[G^{\text{PDIS}}(\tau^{\mu_{0:T-1}}_{0:T-1})]=J(\pi).$
@@ 4 Variance Reduction in Reinforcement Learning
$u_{\pi,t}(s,a)$ is non-negative because it is the sum of three non-negative terms. Therefore, $\sqrt{u_{\pi,t}(s,a)}$ is always well-defined. In (22), $\mu_{t}^{*}(a|s)\propto\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}$ means
$\textstyle\mu_{t}^{*}(a|s)\doteq\frac{\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}}{\sum_{b}\pi_{t}(b|s)\sqrt{u_{\pi,t}(s,b)}}.$ (24)
If $\forall a,\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}=0$, the denominator is zero. In this case, we use the convention to interpret $\mu_{t}^{*}(a|s)$ as a uniform distribution, i.e., $\forall a,\mu_{t}^{*}(a|s)=1/{|\mathcal{A}|}$. We adopt this convention for $\propto$ in the rest of the paper to simplify the presentation. Now, we define the enlarged space $\Lambda$ as
@@ 6 Learning Closed-Form Behavior Policies
Its proof is in Appendix A.8. This lemma allows us to learn $u$ recursively without approximating the complicated trajectory variance. Subsequently, the desired optimal behavior policy $\mu^{*}$ can be easily calculated using (22). To ensure broad applicability, we utilize the behavior policy-agnostic offline learning setting (Nachum et al.,, 2019), in which the offline data consists of $\quantity{(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})}_{i=1}^{m}$, with $m$ previously logged data tuples. Those tuples can be generated by one or multiple behavior policies, regardless of whether these policies are known or unknown, and they are not required to form a complete trajectory. In the $i$-th data tuple, $t_{i}$ represents the time step, $s_{i}$ is the state at time step $t_{i}$, $a_{i}$ is the action executed, $r_{i}$ is the sampled reward, and $s_{i}^{\prime}$ is the successor state. In this paper, we learn $(\mu^{*},b^{*})$ from cheaply available offline data using Fitted $Q$Evaluation (FQE, (Le et al.,, 2019)), but our framework is ready to integrate any state-of-the-art offline policy evaluation technique.
@@ 7 Empirical Results
In this section, we show the empirical results comparing our methods against three baselines: (1) the on-policy Monte Carlo estimator, (2) the offline data informed estimator (ODI, Liu and Zhang, (2024)), and (3) the doubly robust estimator (DR, Jiang and Li, (2016)). In the doubly robust estimator, because they do not design a tailored behavior policy, we leverage the conventional wisdom to use the target policy $\pi$ as the behavior policy. Given previously logged offline data, our method learns the optimal behavior policy and the optimal baseline tuple $(\mu^{*},b^{*})$ using Algorithm 1. All related methods learn their required quantities from the same offline dataset to ensure fair comparisons. Then, we use the behavior policy $\mu^{*}$ for data collection and the baseline $b^{*}$ for data processing. Because our method reduces variance in both the data-collecting and the data-processing phases, we name our method doubly optimal (DOpt) policy evaluation. Our implementation is made publicly available to facilitate future research [^1] Link available upon publication..
@@
Gridworld: We first conduct experiments in Gridworld with $n^{3}$ states, i.e., an $n\times n$ grid with the time horizon also being $n$. The number of states in this Gridworld environment scales cubically with $n$, offering a suitable tool to test algorithm scalability. We choose Gridworld with $n^{3}=1,000$ and $n^{3}=27,000$, which are the largest Gridworld environment tested among related works (Jiang and Li,, 2016; Hanna et al.,, 2017; Liu and Zhang,, 2024). We use randomly generated reward functions with $30$ randomly generated target policies. The offline data is generated by various unknown policies to simulate cheaply available but segmented offline data. Because MC methods use each episode as one empirical return sample, we view each episode as one online sample. We report the relative error of the four methods against the number of online samples. This relative error is the estimation error normalized by the estimation error of the on-policy Monte Carlo estimator after the first episode. Thus, the relative error of the on-policy Monte Carlo estimator starts from $1$.
Figure 1 shows our method outperforms all baselines by a large margin. The blue line in the graph is below all other lines, indicating that our method requires fewer samples to achieve the same accuracy. This is because our designed $(\mu^{*},b^{*})$ substantially reduces estimation variance. In Table 1, we quantify such variance reduction, showing our method reduces variance by around $75\%$ in both Gridworld with size $1,000$ and $27,000$. One observation is that DR performs slightly better than ODI in smaller Gridworld but is slightly worse in larger Gridworld, which shows that there might be no dominating relationship between those two methods. Meanwhile, our method is superior to both approaches because the variance reduction of our method comes from both data-collecting and data-processing. More experiment details are in Appendix B.
Figure 2: Results on MuJoCo. Each curve is averaged over 900 runs (30 target policies, each having 30 independent runs). Shaded regions denote standard errors and are invisible for some curves because they are too small.
@@
Table 2: Episodes needed to achieve the same of estimation accuracy that on-policy Monte Carlo achieves with $1000$ episodes. Standard errors are plotted in Figure 2.
@@ Appendix B Experiment Details
In this paper, we first learn the action-value function $q_{\pi,t}$ from offline data using Fitted Q-Evaluation algorithms (FQE, Le et al., (2019)), but our framework is ready to integrate state-of-the-art offline policy evaluation techniques. Notably, Fitted Q-Evaluation (FQE, Le et al., (2019)) is a different algorithm from Fitted Q-Improvement (FQI). Fitted Q-Evaluation is not prone to overestimate the action-value function $q_{\pi,t}$ because Fitted Q-Evaluation does not have any $\max$ operator and does not change the policy.
@@
the first term is an expectation of $S_{t+1}$. Because we have $(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})$ data tuples, we construct $\nu$ using $s_{i}^{\prime}$ in $(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})$ data tuples as the sample of the first term and compute the rest quantity using the learned action-value function $q_{\pi,t}$ and reward data $r_{i}$. Therefore, we construct $\mathcal{D}_{\nu}\doteq\quantity{(t_{i},s_{i},a_{i},\nu_{i},s^{\prime}_{i})}_{i=1}^{m}$. Finally, by passing data tuples in $\mathcal{D}_{\nu}$ from $t=T-1$ to 00, we fit the function $u_{\pi,t}$ using FQE in a dynamic programming way with respect to the recursive form of $u_{\pi,t}$ derived in Lemma 3. For each time step, we take a copy of the neural network as an approximation to function $u_{\pi,t}$ at time step $t$. After learned function $u_{\pi,t}$ and $q_{\pi,t}$, we return the learned behavior policy $\mu^{*}_{t}(a|s)\propto\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}$ and the learned baseline function $b^{*}_{t}(s,a)=q_{\pi,t}(s,a)$. The pseudocode of this procedure is presented in Algorithm 1.
Figure 3: MuJoCo robot simulation tasks (Todorov et al.,, 2012). Environments from the left to the right are Ant, Hopper, InvertedDoublePendulum, InvertedPendulum, and Walker.

### v2

@@ 1 Introduction
Reinforcement learning (RL, Sutton and Barto, (2018)) has achieved remarkable success in various sequential decision-making problems. For example, RL algorithms have reduced energy consumption for Google data centers’ cooling by $40\%$ (Chervonyi et al.,, 2022), predicted protein structures with competitive accuracy (Jumper et al.,, 2021), and discovered faster matrix multiplication algorithms (Fawzi et al.,, 2022). When applying RL algorithms, policy evaluation plays a critical role to allow practitioners to estimate the performance of a policy before committing to its full deployment and test different algorithmic choices. A commonly used approach among RL practitioners for policy evaluation is the on-policy Monte Carlo method, where a policy (i.e., the target policy) is evaluated by directly executing itself. However, using the target policy itself as the behavior policy is not optimal (Hanna et al.,, 2017; Liu and Zhang,, 2024; Liu et al., 2024b, ; Chen et al.,, 2024), leading to evaluation with high variance. This suboptimality of on-policy evaluation results in extensive needs for collecting online samples to achieve a desired level of accuracy.
@@ 2 Related Work
Reducing the variance for policy evaluation in reinforcement learning (RL) has been widely studied. One rising approach is variance reduction by designing a proper data-collecting policy, also known as the behavior policy. Noticing that the target policy itself is not the best behavior policy, Hanna et al., (2017) formulate the task of finding a variance-reduction behavior policy as an optimization problem. They use stochastic gradient descent to update a parameterized behavior policy. However, the stochastic method has been known to easily get stuck in highly suboptimal points in just moderately complex environments, where various local optimal points exist (Williams,, 1992). By contrast, our method directly learns the globally optimal behavior policy without doing a policy search. Moreover, their method requires highly sensitive hyperparameter tuning to learn the behavior policy effectively. Specifically, the learning rate can vary by up to $10^{5}$ times across different environments, as reported in the experiments of Hanna et al., (2017). This extreme sensitivity requires online tuning, consuming massive online data. By contrast, we propose an efficient algorithm to learn our optimal behavior policy with purely offline data. Furthermore, Hanna et al., (2017) constrain the online data to be complete trajectories. By contrast, our method copes well with incomplete offline data tuples, which is widely applicable.
@@ 3 Background
In this paper, we study a finite horizon Markov Decision Process (MDP, Puterman, (2014)). In this MDP, there is a finite action space $\mathcal{A}$, a finite action space $\mathcal{A}$, a transition probability function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, an initial state distribution $p_{0}:\mathcal{S}\to[0,1]$, and a constant horizon length $T$. For simplifying notations, we consider the undiscounted setting without loss of generality. Our method naturally applies to the discounted setting as long as the horizon is fixed and finite (Puterman,, 2014). We define a shorthand $[n]\doteq\quantity{0,1,\dots,n}$ for any integer $n$.
@@
The MDP process begins at time step 00, where an initial state $S_{0}$ is sampled from $p_{0}$. At each time step $t\in[T-1]$, an action $A_{t}$ is sampled based on $\pi_{t}(\cdot\mid S_{t})$. Here, $\pi_{t}:\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the policy at time step $t$. Then, a finite reward $R_{t+1}\doteq r(S_{t},A_{t})$ is given by the environment and a successor state $S_{t+1}$ is obtained based on $p(\cdot\mid S_{t},A_{t})$. We use abbreviations $\pi_{i:j}\doteq\quantity{\pi_{i},\pi_{i+1},\dots,\pi_{j}}$ and $\pi\doteq\pi_{0:T-1}$. At each time step $t$, the return is defined as $G_{t}\doteq\sum_{i={t+1}}^{T}R_{i}.$ Then, we define the state-value and action-value functions as $v_{\pi,t}(s)\doteq\mathbb{E}_{\pi}\left[G_{t}\mid S_{t}=s\right]$ and $q_{\pi,t}(s,a)\doteq\mathbb{E}_{\pi}\left[G_{t}\mid S_{t}=s,A_{t}=a\right].$ We adopt the total rewards performance metric (Puterman,, 2014) for the measurement of the performance of policy $\pi$, which is defined as $J(\pi)\doteq\sum_{s}p_{0}(s)v_{\pi,0}(s)$. In this work, we use Monte Carlo methods, as introduced by Kakutani, (1945), for estimating the total rewards $J(\pi)$. The most straightforward Monte Carlo method is to draw samples of $J(\pi)$ through the online execution of the policy $\pi$. The empirical average of the sampled returns converges to $J(\pi)$ as the number of samples increases. Since this method estimates a policy by executing itself, it is called on-policy learning (Sutton, 1988).
@@
Moving forward, we focus on off-policy evaluation, where the goal is to estimate the total rewards $J(\pi)$ of an interested policy $\pi$, which is called the target policy. Data for off-policy evaluation are collected by executing a different policy $\mu$, called the behavior policy. In off-policy evaluation, we generate each trajectory $\quantity{S_{0},A_{0},R_{1},S_{1},A_{1},R_{2},\dots,S_{T-1},A_{T-1},R_{T}}$ by a behavior policy $\mu$ with $A_{t}\sim\mu_{t}(\cdot|S_{t}).$ We use a shorthand $\tau^{\mu_{t:T-1}}_{t:T-1}\doteq\quantity{S_{t},A_{t},R_{t+1},\dots,S_{T-1},A_{T-1},R_{T}}$ for a trajectory generated by the behavior policy $\mu$ from time step $t$ to $T-1$ inclusively. We use the importance sampling ratio to reweight the rewards obtained by the behavior policy $\mu$, in order to give an estimate of $J(\pi)$. We define the importance sampling ratio at time step $t$ as $\rho_{t}\doteq\frac{\pi_{t}(A_{t}\mid S_{t})}{\mu_{t}(A_{t}\mid S_{t})}.$ Then, the product of importance sampling ratios from time $t$ to $t^{\prime}\geq t$ is defined as $\rho_{t:t^{\prime}}\doteq\prod_{k=t}^{t^{\prime}}\frac{\pi_{k}(A_{k}|S_{k})}{\mu_{k}(A_{k}|S_{k})}.$ In off-policy learning, there are several ways to use the importance sampling ratios (Geweke,, 1988; Hesterberg,, 1995; Koller and Friedman,, 2009; Thomas,, 2015). In this paper, we investigate the per-decision importance sampling estimator (PDIS, Precup et al., (2000)) and leave the investigation of others for future work. We define the PDIS Monte Carlo estimator as $G^{\text{PDIS}}(\tau^{\mu_{t:T-1}}_{t:T-1})\doteq\sum_{k=t}^{T-1}\rho_{t:k}R_{k+1},$ which can also be expressed recursively as
@@
Under the classic policy coverage assumption (Precup et al.,, 2000; Maei,, 2011; Sutton et al.,, 2016; Zhang,, 2022; Liu et al., 2024a, ) $\forall t,s,a,\mu_{t}(a|s)=0\implies\pi_{t}(a|s)=0$, this off-policy estimator $G^{\text{PDIS}}(\tau^{\mu_{0:T-1}}_{0:T-1})$ provides an unbiased estimation for $J(\pi)$, i.e., $\mathbb{E}\quantity[G^{\text{PDIS}}(\tau^{\mu_{0:T-1}}_{0:T-1})]=J(\pi).$
@@ 4 Variance Reduction in Reinforcement Learning
$u_{\pi,t}(s,a)$ is non-negative because it is the sum of three non-negative terms. Therefore, $\sqrt{u_{\pi,t}(s,a)}$ is always well-defined. In (22), $\mu_{t}^{*}(a|s)\propto\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}$ means $\mu_{t}^{*}(a|s)\doteq\frac{\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}}{\sum_{b}\pi_{t}(b|s)\sqrt{u_{\pi,t}(s,b)}}$. If $\forall a,\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}=0$, the denominator is zero. In this case, we use the convention to interpret $\mu_{t}^{*}(a|s)$ as a uniform distribution, i.e., $\forall a,\mu_{t}^{*}(a|s)=1/{|\mathcal{A}|}$. We adopt this convention for $\propto$ in the rest of the paper to simplify the presentation. Now, we define the enlarged space $\Lambda$ as
@@ 6 Learning Closed-Form Behavior Policies
Its proof is in Appendix A.8. This lemma allows us to learn $u$ recursively without approximating the complicated trajectory variance. Subsequently, the desired optimal behavior policy $\mu^{*}$ can be easily calculated using (22). To ensure broad applicability, we utilize the behavior policy-agnostic offline learning setting (Nachum et al.,, 2019), in which the offline data consists of $\quantity{(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})}_{i=1}^{m}$, with $m$ previously logged data tuples. Those tuples can be generated by various unknown behavior policies, and they are not required to form a complete trajectory. In the $i$-th data tuple, $t_{i}$ represents the time step, $s_{i}$ is the state at time step $t_{i}$, $a_{i}$ is the action executed, $r_{i}$ is the sampled reward, and $s_{i}^{\prime}$ is the successor state. In this paper, we learn $(\mu^{*},b^{*})$ from cheaply available offline data using Fitted $Q$Evaluation (FQE, (Le et al.,, 2019)), but our framework is ready to integrate any state-of-the-art offline policy evaluation technique. As for constructing $\nu$, we use the learned $q$ function and $r_{i}$, $s^{\prime}_{i}$ from the data tuples, according to the derivation (232) in Appendix B.
@@ 7 Empirical Results
In this section, we show the empirical comparison between our methods and three baselines: (1) the on-policy Monte Carlo estimator, (2) the offline data informed estimator (ODI, Liu and Zhang, (2024)), and (3) the doubly robust estimator (DR, Jiang and Li, (2016)). In the doubly robust estimator, because they do not design a tailored behavior policy, we leverage the conventional wisdom to use the target policy $\pi$ as the behavior policy. Given previously logged offline data, we learn our optimal behavior policy and the optimal baseline tuple $(\mu^{*},b^{*})$ using Algorithm 1. All baseline methods learn their required quantities from the same offline dataset to ensure fair comparisons. We use the behavior policy $\mu^{*}$ for data collection and the baseline $b^{*}$ for data processing. Since our method reduces variance in both the data-collecting and the data-processing phases, we name our method doubly optimal (DOpt) policy evaluation. More experiment details are in Appendix B.
@@
Gridworld: We begin by conducting experiments in Gridworld with $n^{3}$ states, i.e., an $n\times n$ grid with $n$ as the time horizon. The number of states in this Gridworld environment scales cubically with $n$, offering a suitable tool to test algorithm scalability. We choose Gridworld with $n^{3}=1,000$ and $n^{3}=27,000$, which are the largest Gridworld environment tested among related works (Jiang and Li,, 2016; Hanna et al.,, 2017; Liu and Zhang,, 2024). We use randomly generated reward functions with $30$ randomly generated target policies. The offline data is generated by various unknown policies to simulate cheaply available but segmented offline data. Because MC methods use each episode as one empirical return sample, we view each episode as one online sample. We report the relative error of the four methods against the number of online samples. This relative error is the estimation error normalized by the estimation error of the on-policy Monte Carlo estimator after the first episode. Thus, the relative error of the on-policy Monte Carlo estimator starts from $1$.
Figure 1 shows our method outperforms all baselines by a large margin. The blue line in the graph is below all other lines, indicating that our method requires fewer samples to achieve the same accuracy. This is because our designed $(\mu^{*},b^{*})$ substantially reduces estimation variance. In Table 1, we quantify such variance reduction, showing our method reduces variance by around $75\%$ in both Gridworld with size $1,000$ and $27,000$.
One observation is that DR performs slightly better than ODI in smaller Gridworld but is slightly worse in larger Gridworld, which shows that there might be no dominating relationship between those two methods. Meanwhile, our method is superior to both approaches because the variance reduction of our method comes from both data-collecting and data-processing.
Figure 2: Results on MuJoCo. Each curve is averaged over 900 independent runs (30 target policies, each having 30 independent runs). Shaded regions denote standard errors and are invisible for some curves because they are too small.
@@
Table 2: Episodes needed to achieve the same of estimation accuracy that on-policy Monte Carlo achieves with $1000$ episodes. Standard errors are plotted in Figure 2. Each number is averaged over $900$ independet runs.
@@ Appendix B Experiment Details
In this paper, we first learn the action-value function $q_{\pi,t}$ from offline data using Fitted Q-Evaluation algorithms (FQE, Le et al., (2019)), but our method can integrate state-of-the-art offline policy evaluation techniques. Notably, Fitted Q-Evaluation (FQE, Le et al., (2019)) is a different algorithm from Fitted Q-Improvement (FQI). Fitted Q-Evaluation is not prone to overestimate the action-value function $q_{\pi,t}$ because Fitted Q-Evaluation does not have any $\max$ operator and does not change the policy.
@@
the first term is an expectation of $S_{t+1}$. Because we have $(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})$ data tuples, we construct $\nu$ using $s_{i}^{\prime}$ in $(t_{i},s_{i},a_{i},r_{i},s_{i}^{\prime})$ data tuples as the sample of the first term and compute the rest quantity using the learned action-value function $q_{\pi,t}$ and reward data $r_{i}$. Therefore, we construct $\mathcal{D}_{\nu}\doteq\quantity{(t_{i},s_{i},a_{i},\nu_{i},s^{\prime}_{i})}_{i=1}^{m}$. Finally, by passing data tuples in $\mathcal{D}_{\nu}$ from $t=T-1$ to 00, we fit the function $u_{\pi,t}$ using FQE in a dynamic programming way with respect to the recursive form of $u_{\pi,t}$ derived in Lemma 3. For each time step, we take a copy of the neural network as an approximation of function $u_{\pi,t}$ at time step $t$. After learning the functions $u_{\pi,t}$ and $q_{\pi,t}$, we return the learned behavior policy $\mu^{*}_{t}(a|s)\propto\pi_{t}(a|s)\sqrt{u_{\pi,t}(s,a)}$ and the learned baseline function $b^{*}_{t}(s,a)=q_{\pi,t}(s,a)$. The pseudocode of this procedure is presented in Algorithm 1.
Figure 3: MuJoCo robot simulation tasks (Todorov et al.,, 2012). The pictures are adapted from (Liu and Zhang,, 2024). Environments from the left to the right are Ant, Hopper, InvertedDoublePendulum, InvertedPendulum, and Walker.